OpenStudy (anonymous):

How to find the x and y components of the following vectors? a. Magnitude = 9.30m; 0=60° b. Magnitude = 22.0km; 0= 135° c. Magnitude = 6.35cm ; 0 = 307°

OpenStudy (aaronq):

Use trigonometry  \(\sf sin\theta=\dfrac{opposite}{hypotenuse}\) and \(\sf cos\theta=\dfrac{adjacent}{hypotenuse}\)